# A method to benchmark high-dimensional process drift detection

### v1

@@ 4. The temporal area under the curve
In this section, we construct a score to measure the predictive power of machine learning models for process drift detection. In order to do so, we first formalize what we understand as a process drift and which assumptions we require. Let $C_{1},\ldots,C_{T}$ be a sequence of process curves and let $\mathcal{D}\subset[T]$ be the set of curve indices belonging to drifts. Our first assumption is that drifts, different than point anomalies, appear sequentially and can be uniquely decomposed into disjoint segments:
@@
The drift segments can be considered as a partition of the smallest consecutive drifts which cannot decomposed any further into smaller segments. Now, assume we also have the output $s\in\mathbb{R}^{T}$ of a detector where each coordinate $s_{t}$ quantifies how likely the curve $C_{t}$ of the $i$-t-h process execution belongs to a drift, that is, the higher $s_{t}$ the more likely the detector classifies $t\in{\mathcal{D}}$. By choosing a threshold $\tau\in\mathbb{R}$, we can construct a set
@@
which serves as a possible candidate for ${\mathcal{D}}$. Clearly, if $\tau_{1}\geq\tau_{2}$, then $\hat{\mathcal{D}}(s,\tau_{1})\subseteq\hat{\mathcal{D}}(s,\tau_{2})$. Its also straight-forward to see that for every $\tau$, the set $\hat{\mathcal{D}}(s,\tau)$ decomposes uniquely into drift segments $\hat{\mathcal{D}}_{1},\ldots,\hat{\mathcal{D}}_{l}$ as defined in Definition 4.1 and that the length and number of these atomic segments depends on $\tau$. Now, to quantify the predictive power of the detector yielding $s$, one needs to quantify how close $\hat{\mathcal{D}}(s,\cdot)$ is to ${\mathcal{D}}$ when $\tau$ varies. There are many established set-theoretic measurements that are widely used in practice to quantify the distance between two finite and binary sets $A$ and $B$, like the Jaccard index $\frac{|A\cap B|}{|A\cup B|}$, the Hamming distance $|A\setminus B|+|B\setminus A|$, or the Overlap coefficient $\frac{|A\cap B|}{\min(|A|,|B|)}$ just to name a few. Most scores, however, have as a build-in assumption that the elements of the set are iid and hence the temporal context is largely ignored making them unsuitable for process drift detection. Moreover, for most detectors we have to select a discrimination threshold $\tau$, making evaluation cumbersome as it requires to tune the threshold on a separate held-out dataset. Moreover, in most practical scenarios, ${\mathcal{D}}$ is only a small subset and thus the evaluation metric has to consider highly imbalanced szenarios as well.
@@
Clearly, detectors are required where all true drift segments ${\mathcal{D}}_{i}$ are overlapped by predicted drift segments. For this, let $T_{i}:=\{j\in[l]:{\mathcal{D}}_{i}\cap\hat{\mathcal{D}}_{j}\neq\emptyset\}$. Clearly, $T_{i}\cap T_{i+1}\neq\emptyset$ if ${\mathcal{D}}_{i}$ and ${\mathcal{D}}_{i+1}$ both intersect with a predicted drift segment. Now, the set ${\mathcal{T}}_{i}:=\cup_{j\in T_{i}}\hat{\mathcal{D}}_{j}$ which is the union of all predictive segments intersecting with ${\mathcal{D}}_{i}$ serves as a candidate for ${\mathcal{D}}_{i}$. To measure how well ${\mathcal{D}}_{i}$ is covered - or overlapped - by ${\mathcal{T}}_{i}$ we define the soft overlap score inspired by the Overlap coefficient as follows:
@@
Obviously, an sOLS of $1$ is best possible and this is reached if and only if ${\mathcal{T}}_{i}={\mathcal{D}}_{i}$. It is easy to see that for fixed ${\mathcal{D}}_{i}$, the enlargement of ${\mathcal{T}}_{i}$ beyond the boundaries of ${\mathcal{D}}_{i}$ improves the overlap score, as $|{\mathcal{T}}_{i}|$ increases and one of either $\max({\mathcal{T}}_{i}\cup{\mathcal{D}}_{i})$ or $-\min(T_{i}\cup{\mathcal{D}}_{i})$ increases as well. A special case is if ${\mathcal{D}}_{i}$ is completely covered by ${\mathcal{T}}_{i}$, i.e. ${\mathcal{D}}_{i}\subseteq{\mathcal{T}}_{i}$, then it follows that ${\mathcal{T}}_{i}$ is an interval as well and thus $\text{sOLS}({\mathcal{D}}_{i},s,\tau)=1$. When ${\mathcal{T}}_{i}$ enlarges, then the number of false positives, i.e. the time points $t$ contained in some $\hat{\mathcal{D}}_{i}$ and in the complement $\overline{{\mathcal{D}}}:=[T]\setminus{\mathcal{D}}$ of the ground truth ${\mathcal{D}}$, enlarges as well. Thus, the predictive power of a detector is shown in the overlap score as well as the created false positive rate
@@
Output: Overlap score
@@
3:$s\leftarrow\mathbf{0}\in\mathbb{R}^{k}$
@@
5:     $T_{i}\leftarrow\{j\in[l]:\hat{\mathcal{D}}_{j}\cap{\mathcal{D}}_{i}\neq\emptyset\}$triangleright All predicted drift segments overlapping with ${\mathcal{D}}_{i}$
6:     $\mathcal{T}_{i}\leftarrow\cup_{j\in T_{i}}\hat{\mathcal{D}}_{j}$triangleright Union of all segments intersecting with ${\mathcal{D}}_{i}$
7:     $s_{i}\leftarrow\frac{|\mathcal{T}_{i}\cap{\mathcal{D}}_{i}|}{\max(\mathcal{T}_{i}\cup\mathcal{D}_{i})-\min(\mathcal{T}_{i}\cup\mathcal{D}_{i})+1}$triangleright fraction of overlap
@@
9:return $\frac{1}{k}\sum^{k}_{i=1}s_{i}$
Our score considers both, the OLS as well as the FPR, which mutually influence each other. In the computation of the AUC, any threshold $\tau$ from $[\min_{t}(s_{t}),\max_{t}(s_{t})]$ yields a pair of false positive rate $\textnormal{FPR}({\mathcal{D}},s,\tau)$ and true positive rate $\textnormal{TPR}({\mathcal{D}},s,\tau)$ which can be drawn as a curve in the space where FPR is on the $x$-axis and TPR on the $y$-axis. Similarly, we define the temporal area under the curve, or just TAUC, as the area under the FPR-OLS curve while the discrimination threshold $\tau$ varies. We refer to the soft TAUC, or just sTAUC, to the area under the FPR-sOLS curve (see Figure 4).
Note that the integral of the curve can be computed using two different methods, the step rule and the trapezoidal rule and depending on which method is used, the value of the score may differs. We showcase this behavior in detail for trivial detectors in Appendix C. In Appendix B, we investigate in several synthetic cases in depth the differences and similarities between sTAUC, TAUC, and AUC.
@@
Consider the situation shown in Figure 3. There, we have two true drift segments $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and three segments $\hat{\mathcal{D}}_{1}$, $\hat{\mathcal{D}}_{2}$ and $\hat{\mathcal{D}}_{3}$ as drift segments of some detector output $\hat{\mathcal{D}}(s,\tau)$. Clearly, $T_{1}=\{1,2\}$ where $T_{2}=\{3\}$ as only ${\mathcal{D}}_{3}$ overlaps with ${\mathcal{D}}_{2}$. To unveil ${\mathcal{D}}_{1}$, the detector needs to separate drift segments, leading to false negatives and positives and thus a relatively low OLS. On the other hand, as $\mathcal{\hat{D}}_{3}\subset\mathcal{D}_{2}$, we have $\mathcal{T}_{2}=\mathcal{\hat{D}}_{3}$.
@@ 5. Experiments
Next, we benchmark existing algorithms on data generated with our framework driftbench and reporting the TAUC [^1] All datasets and algorithms used are available in the repository of driftbench.. The goal of the benchmark is to provide a proof of concept for our score and data generation method, not to be very comprehensive on the model side. Thus, based on our literature research in Section 1 we have hand-selected a small set of typically used model patterns drift detectors used in practice consists of (see Section 5.1).
The basic evaluation loop follows a typical situation from manufacturing, where process engineers have to identify time periods within a larger curve datasets where the process has drifted. Thus, all models consume as input a process curve dataset $C_{1},\ldots,C_{T}$ and do not have access to the ground truth ${\mathcal{D}}$, which is the set of curves belonging to a drift (see Section 5.3). Afterwards, each model predicts for each curve $C_{t}$ from this dataset, a score $s_{t}$ on which then the TAUC, sTAUC, and AUC are computed. To account for robustness, we generate each dataset of a predefined specification multiple times with different random seeds each, leading to slightly different datasets of roughly same complexity.
@@ 5.1. Algorithms
The algorithms used can be decomposed into multiple steps (see also Figure 5), but not all algorithms use all steps. First, there are may some features extracted from each curve. Afterwards, a sliding window collects and may aggregates these such that a score is computed.
@@ 5.1.1. Feature extraction
In this steps, we use autoencoders [25] to compute a $k$ dimensional representation $e_{t}\in\mathbb{R}^{k}$ for each high-dimensional process curve $C_{t}$ with $k$ small. The indention behind is to estimate an inverse of the unknown function $f$ and to recover information about the support points used. Moreover, we also apply deterministic aggregations over the $x$-information of each curve $C_{t}$,
@@ 5.1.2. Windowing and aggregation
In this step, the algorithms may aggregate the data from the previous step using a fixed window of size $m$ that is applied in a rolling fashion along the process iterations. One aggregation we use is to first compute for each coordinate $j\in[k]$ of $e_{t}\in\mathbb{R}^{k}$ with $t\geq m$ the rolling mean:
(5.1) $a_{t,j}=\frac{1}{m}\sum^{t}_{i=t-m+1}e_{i,j}.$
These values can then further be statistically aggregated, like by taking the maximum $a_{t}:=\max\{a_{t,j}:j\in[k]\}$.
@@ 5.1.3. Score computing
Another approach follows a probabilistic path by testing if a set of subsequent datapoints $\{a_{t-m+1},\ldots,a_{t}\}$ come from the same distribution as a given reference set. In our study, we use a windowed version [23] of the popular Kolmogorov-Smirnov test [26], often called KSWIN, which makes no assumption of the underlying data distribution. However, this can only be applied when $a_{t}$ is a scalar. More particularly, we define two window sizes, $m_{r}$ for the reference data and $m_{o}$ for the observation. The windows are offset by constant $\delta>0$. We then invoke the KS-test and receive a $p$-value $p_{t}$, which is small if the datasets come from different distributions. Thus, one way to derive a final score is to compute $s_{t}=\log(1+\frac{1}{p_{t}})$. We also evaluate algorithms that derive their score based on a similarity search within $\{a_{1},\ldots,a_{t}\}$. Here, we use clustering algorithms, like the popular $k$-means algorithm, and use the euclidean distance to the computed cluster center of $a_{t}$ as $s_{t}$. Another way is to fit a probability density function on $s_{t}$, like a mixture of Gaussian distributions, and to set $s_{t}$ as the log likelihood of $a_{t}$ within this model.
@@ 5.2. Algorithm Overview
  • Cluster($n_{c}$): A cluster algorithm performed on the raw curves using $n_{c}$ clusters where score is distance to closest cluster center.
@@ 5.3. Datasets
Fig. 6. The datasets used in our benchmark study. The true drift segments are marked in green. Lower figures show selected curves, whose color encodes the process iteration $t\in[T]$ blue marks smaller $t$ values, red larger ones.
To generate dataset-1 and dataset-2, we used
$f(w,x)=\sum_{i=0}^{7}w_{i}\cdot x^{i}$
@@
$f(w,x)=w_{0}\cdot x\cdot\sin(\pi\cdot x-w_{1})+w_{2}\cdot x$
and generated $T=10.000$ many curves, each having $|I|=100$ datapoints. It only holds a single drift, where the global minimum at drifts consistently over a small period of time along the $x$-axis. In all datasets, the relative number of curves belonging to a drift is very small: roughly $2$ percent in dataset-1, $0.1$ percent in dataset-2 and $1$ percent in dataset-3.

### v2

@@ 4. The temporal area under the curve
Different usecases require different performance metrics to evaluate algorithms. In classification, for instance, sometimes avoiding false positives is, sometimes avoiding false negatives. However, when it comes to general benchmarking classifiers somewhat independently of their precise application in the sense to see how well their response correlates to the actual class label, the AUC [26] is frequently used. However, the vanilla AUC takes samples independently of their temporal context, that is, independent of samples from the previous and next process execution. Thus, we construct in this section a more suitable metric to measure the predictive power of machine learning models for process drift detection. In order to do so, we first formalize what we understand as a process drift and which assumptions we require. Let $C_{1},\ldots,C_{T}$ be a sequence of process curves and let $\mathcal{D}\subset[T]$ be the set of curve indices belonging to drifts. Our first assumption is that drifts, different from point anomalies, appear sequentially and can be uniquely decomposed into disjoint segments:
@@
Refer to captionRefer to caption
Fig. 5. Applying a process drift detector on each process curves yields a score $s$ which needs to be compared to the ground truth $\mathcal{D}$ for each threshold $\tau$.
The drift segments can be considered as a partition of the smallest consecutive drifts which cannot decomposed any further into smaller segments. Now, assume we also have the output $s\in\mathbb{R}^{T}$ of a detector where each coordinate $s_{t}$ quantifies how likely the curve $C_{t}$ of the $i$-t-h process execution belongs to a drift, that is, the higher $s_{t}$ the more likely the detector classifies $t\in{\mathcal{D}}$ (see also Figure 5). By choosing a threshold $\tau\in\mathbb{R}$, we can construct a set
@@
which serves as a possible candidate for ${\mathcal{D}}$. Clearly, if $\tau_{1}\geq\tau_{2}$, then $\hat{\mathcal{D}}(s,\tau_{1})\subseteq\hat{\mathcal{D}}(s,\tau_{2})$. Its also straight-forward to see that for every $\tau$, the set $\hat{\mathcal{D}}(s,\tau)$ decomposes uniquely into drift segments $\hat{\mathcal{D}}_{1},\ldots,\hat{\mathcal{D}}_{l}$ as defined in Definition 4.1 and that the length and number of these atomic segments depends on $\tau$. Now, to quantify the predictive power of the detector yielding $s$, one needs to quantify how close $\hat{\mathcal{D}}(s,\cdot)$ is to ${\mathcal{D}}$ when $\tau$ varies. There are many established set-theoretic measurements that are widely used in practice to quantify the distance between two finite and binary sets $A$ and $B$, like the Jaccard index $\frac{|A\cap B|}{|A\cup B|}$, the Hamming distance $|A\setminus B|+|B\setminus A|$, or the Overlap coefficient $\frac{|A\cap B|}{\min(|A|,|B|)}$ just to name a few. Most metrics, however, have as a build-in assumption that the elements of the set are iid and hence the temporal context is largely ignored making them unsuitable for process drift detection. Moreover, for most detectors we have to select a discrimination threshold $\tau$, making evaluation cumbersome as it requires to tune the threshold on a separate held-out dataset. Moreover, in most practical scenarios, ${\mathcal{D}}$ is only a small subset and thus the evaluation metric has to consider highly imbalanced scenarios as well.
@@
Clearly, detectors are required where all true drift segments ${\mathcal{D}}_{i}$ are overlapped by predicted drift segments. For this, let $L_{i}:=\{j\in[l]:{\mathcal{D}}_{i}\cap\hat{\mathcal{D}}_{j}\neq\emptyset\}$. Clearly, $L_{i}\cap L_{i+1}\neq\emptyset$ if ${\mathcal{D}}_{i}$ and ${\mathcal{D}}_{i+1}$ both intersect with a predicted drift segment. Now, the set ${\mathcal{T}}_{i}:=\cup_{j\in L_{i}}\hat{\mathcal{D}}_{j}$ which is the union of all predictive segments intersecting with ${\mathcal{D}}_{i}$ serves as a candidate for ${\mathcal{D}}_{i}$. To measure how well ${\mathcal{D}}_{i}$ is covered - or overlapped - by ${\mathcal{T}}_{i}$ we define the soft overlap score inspired by the Overlap coefficient as follows:
@@
Obviously, an sOLS of $1$ is the best possible and this is reached if and only if ${\mathcal{T}}_{i}={\mathcal{D}}_{i}$. It is easy to see that for fixed ${\mathcal{D}}_{i}$, the enlargement of ${\mathcal{T}}_{i}$ beyond the boundaries of ${\mathcal{D}}_{i}$ improves the overlap score, as $|{\mathcal{T}}_{i}|$ increases and one of either $\max({\mathcal{T}}_{i}\cup{\mathcal{D}}_{i})$ or $-\min(T_{i}\cup{\mathcal{D}}_{i})$ increases as well. A special case is if ${\mathcal{D}}_{i}$ is completely covered by ${\mathcal{T}}_{i}$, i.e. ${\mathcal{D}}_{i}\subseteq{\mathcal{T}}_{i}$, then it follows that ${\mathcal{T}}_{i}$ is an interval as well and thus $\text{sOLS}({\mathcal{D}}_{i},s,\tau)=1$. When ${\mathcal{T}}_{i}$ enlarges, then the number of false positives, i.e. the time points $t$ contained in some $\hat{\mathcal{D}}_{i}$ and in the complement $\overline{{\mathcal{D}}}:=[T]\setminus{\mathcal{D}}$ of the ground truth ${\mathcal{D}}$, enlarges as well. Thus, the predictive power of a detector is shown in the overlap score as well as the created false positive rate
@@
Output: Overlap score .
@@
3:$o\leftarrow\mathbf{0}\in\mathbb{R}^{k}$
@@
5:     $L_{i}\leftarrow\{j\in[l]:\hat{\mathcal{D}}_{j}\cap{\mathcal{D}}_{i}\neq\emptyset\}$triangleright All predicted drift segments overlapping with ${\mathcal{D}}_{i}$
6:     $\mathcal{T}_{i}\leftarrow\cup_{j\in L_{i}}\hat{\mathcal{D}}_{j}$triangleright Union of all segments intersecting with ${\mathcal{D}}_{i}$
7:     $o_{i}\leftarrow\frac{|\mathcal{T}_{i}\cap{\mathcal{D}}_{i}|}{\max(\mathcal{T}_{i}\cup\mathcal{D}_{i})-\min(\mathcal{T}_{i}\cup\mathcal{D}_{i})+1}$triangleright fraction of overlap
@@
9:return $\frac{1}{k}\sum^{k}_{i=1}o_{i}$
Our score considers both, the OLS and the FPR, which mutually influence each other. In the computation of the AUC, any threshold $\tau$ from $[\min_{t}(s_{t}),\max_{t}(s_{t})]$ yields a pair of false positive rate $\textnormal{FPR}({\mathcal{D}},s,\tau)$ and true positive rate $\textnormal{TPR}({\mathcal{D}},s,\tau)$ which can be drawn as a curve in the space where FPR is on the $x$-axis and TPR on the $y$-axis. Similarly, we define the temporal area under the curve, or just TAUC, as the area under the FPR-OLS curve while the discrimination threshold $\tau$ varies. We refer to the soft TAUC, or just sTAUC, to the area under the FPR-sOLS curve (see Figure 7).
Note that the integral of the curve can be computed using two different methods, the step rule and the trapezoidal rule and depending on which method is used, the value of the score may differ. We showcase this behavior in detail for trivial detectors in Appendix D. In Appendix C, we investigate in several synthetic cases in depth the differences and similarities between sTAUC, TAUC, and AUC.
@@
Consider the situation shown in Figure 6. There, we have two true drift segments $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and three segments $\hat{\mathcal{D}}_{1}$, $\hat{\mathcal{D}}_{2}$ and $\hat{\mathcal{D}}_{3}$ as drift segments of some detector output $\hat{\mathcal{D}}(s,\tau)$. Clearly, $L_{1}=\{1,2\}$ where $L_{2}=\{3\}$ as only ${\mathcal{D}}_{3}$ overlaps with ${\mathcal{D}}_{2}$. To unveil ${\mathcal{D}}_{1}$, the detector needs to separate drift segments, leading to false negatives and positives and thus a relatively small OLS. On the other hand, as $\mathcal{\hat{D}}_{3}\subset\mathcal{D}_{2}$, we have $\mathcal{T}_{2}=\mathcal{\hat{D}}_{3}$.
@@ 5. Experiments
Next, we benchmark existing algorithms on data generated with our framework driftbench and reporting the TAUC. All datasets and algorithms used are available in the repository of driftbench. The goal of the benchmark is to provide a proof of concept for our score and data generation method, not to be very comprehensive on the model side. Thus, based on our literature research in Section 1 we have hand-selected a small set of typically used model patterns drift detectors used in practice consists of (see Section 5.1).
The basic evaluation loop follows a typical situation from manufacturing, where process engineers have to identify time periods within a larger curve datasets where the process has drifted. Thus, all models consume as input a process curve dataset $C_{1},\ldots,C_{T}$ and do not have access to the ground truth ${\mathcal{D}}$, which is the set of curves belonging to a drift (see Section 5.3). Afterwards, each model predicts for each curve $C_{t}$ from this dataset a score $s_{t}\in\mathbb{R}$, and afterwards, the TAUC, sTAUC, and AUC are computed for $s=(s_{1},\ldots,s_{T})$. To account for robustness, we generate each dataset of a predefined specification five times for a different random seed each, leading to slightly different datasets of roughly same complexity. All models are trained unsupervised, i.e. without any information of the true drift segments.
@@ 5.1. Algorithms
The algorithms used can be decomposed into multiple steps (see also Figure 8), but not all algorithms use all steps. First, there are may some features extracted from each curve. Afterwards, a sliding window collects and may aggregate these such that a score is computed.
@@ 5.1.1. Feature extraction
In this step, we use autoencoders [11] to compute a $k$-dimensional representation $e_{t}\in\mathbb{R}^{k}$ for each high-dimensional process curve $C_{t}$ with $k$ small. The indention behind is to estimate an inverse of the unknown function $f$ and to recover information about the support points used. Moreover, we also apply deterministic aggregations over the $x$-information of each curve $C_{t}$.
@@ 5.1.2. Windowing and aggregation
In this step, the algorithms may aggregate the data from the previous step using a fixed window of size $m$ that is applied in a rolling fashion along the process iterations. One aggregation we use is to first compute for each coordinate $j\in[k]$ of $e_{t}\in\mathbb{R}^{k}$ with $t\geq m$ the rolling mean $a_{t,j}=\frac{1}{m}\sum^{t}_{i=t-m+1}e_{i,j}$. These values can then further be statistically aggregated, like by taking the maximum $a_{t}:=\max\{a_{t,j}:j\in[k]\}$.
@@ 5.1.3. Score computing
Another approach follows a probabilistic path by testing if a set of subsequent datapoints $\{a_{t-m+1},\ldots,a_{t}\}$ come from the same distribution as a given reference set. In our study, we use a windowed version [17] of the popular Kolmogorov-Smirnov test [36], often called KSWIN, which makes no assumption of the underlying data distribution. However, this can only be applied when $a_{t}$ is a scalar. More particularly, we define two window sizes, $m_{r}$ for the reference data and $m_{o}$ for the observation. The windows are offset by constant $\delta>0$. We then invoke the KS-test and receive a $p$-value $p_{t}$, which is small if the datasets come from different distributions. Thus, one way to derive a final score is to compute $s_{t}=\log(1+\frac{1}{p_{t}})$. Another probabilistic method we use in our study based on a multivariate statistical test is the Maximum Mean Discrepancy (MMD) [19]. This method uses feature mappings based on kernels, and calculates the distance between the means in these mappings. MMD also makes no assumption about the underlying distribution, and works on multidimensional data. We use this method in the same way using two windows as described in the KS-test. We also evaluate algorithms that derive their score based on a similarity search within $\{a_{1},\ldots,a_{t}\}$. Here, we use clustering algorithms, like the popular $k$-means algorithm, and use the euclidean distance to the computed cluster center of $a_{t}$ as $s_{t}$. Another way is to fit a probability density function on $s_{t}$, like a mixture of Gaussian distributions, and to set $s_{t}$ as the log likelihood of $a_{t}$ within this model.
@@ 5.2. Algorithm Overview
  • Cluster($n_{c}$): A cluster algorithm performed on the raw curves using $n_{c}$ clusters where score is distance to the closest cluster center.
@@
  • AE($k$)-MMD($m_{r},m_{o},\delta$): First, an autoencoder is applied extracting computing $k$ many latent dimensions. Afterwards, a $k$-dimensional sliding MMD-test is applied with two windows of sizes $m_{r}$ and $m_{o}$, where the windows are offset by $\delta$.
@@ 5.3. Datasets
Fig. 9. The datasets used in our benchmark study. The true drift segments are marked in green. Lower figures show selected curves, whose color encodes the process iteration $t\in[T]$ blue marks smaller $t$ values, red larger ones. Recall that dataset-$k$ has $k$ many drift segments.
@@
$\mu(t)=\begin{cases}a,&\textnormal{ if }t<t_{0}\\ b\cdot\frac{t-t_{0}}{t_{1}-t_{0}}+a,&\textnormal{ if }t_{0}\leq t\leq t_{1}\\ b,&\textnormal{ if }t_{1}<t\end{cases}.$